Mohammad Aran a financial planner recommended that 40% of a client cash account be invested in preferred stock earning 9% annual simple interest. The remainder of the client cash was placed in treasury bonds earning 7% annual simple interest. The total annual interest earned from two investment was $2496. Find the total invested
Answer by checkley79(3341)

.09*.40X+.07*.60X=2,496
.036X+.042X=2,496
.078X=2,496
X=2,496/.078
X=$32,000 TOTAL AMOUNT INVESTED.
.09*.40*32,000=$1,152 AMOUNT INVESTED @ 9%.
.07*.60*32,000=$1,344 AMOUNT INVESTED @ 7%.
PROOF:
1152+1344=2496